A diver is 57.5 ft under the surface of a lake (p=1000 kg/m3). What are the absolute and gage pressures at that depth? Answer in both English and SI units.
Answer #1

h = 57.5 ft = 17.526 m

gauge pressure

Pg = h * rho * g = 17.526* 1000* 9.8

Pg = 1.718* 10^5 Pa (SI unit)

Pg = 24.917 psi ( English unit)

=======

absolute pressure

Pa = Patm + Pg

Pa = 1.013 * 10^5 + 1.718* 10^5

Pa = 2.73* 10^5 Pa (SI units)

Pa = 39.595 psi (English unit)

========

Comment before rate in case any doubt, will reply for sure.. goodluck